OpenStudy (linn99123):

what is the median of 60,20,90,100,140,30,80,70

OpenStudy (linn99123):

@undeadknight26 @TheRealMeeeee

OpenStudy (texaschic101):

if there is an even number of numbers, then there will be two middle numbers. In that case, you add them, then divide by 2. Put the numbers in order... 20,30,60,70,80,90,100,140 middle numbers are 70 and 80 (70 + 80) / 2 = 150/2 = 75...scorpion is right :)

OpenStudy (anonymous):

The answer is 75

OpenStudy (anonymous):

do u still have a question on how to do this?